The solution of | (x - 1) + 2 | = 1 is
A. 1 B. 2 C. 5 D. f

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
\left | x-1+2 \right |=1
\left | x+1 \right |=1
x+1=1 or x+1=-1
so x=0 or x=-2
Thanks and Regards
Shaik Aasif
askIITians faculty